OpenStudy (anonymous):

Implicit differentation -> dy/dx y(root(x^2+y^2))=15

OpenStudy (anonymous):

\[y \sqrt{x^2+y^2}=15\]

OpenStudy (anonymous):

calculate y'(x)

OpenStudy (anonymous):

use the product rule first...

OpenStudy (anonymous):

\[y ' \sqrt{x^2+y^2} +\frac{ y }{ 2 \sqrt{x^2+y^2} }*(2x+2y*y')\]

OpenStudy (anonymous):

\[y ' \sqrt{x^2+y^2} +\frac{ 2xy +2y^2y'}{ 2 \sqrt{x^2+y^2} } =0\]

OpenStudy (anonymous):

can you take it from here?

OpenStudy (anonymous):

i want to say yes, i just dont want to mess it up any more. ive been working on this one for soooo long. seems that i get close to the answer but its always off.

OpenStudy (anonymous):

so you did the differentiation right and now you're having algebra problems?

OpenStudy (anonymous):

im sorry about this, ive been just cramming for this test and i think im a little burnt...

OpenStudy (anonymous):

i hate to say it @Algebraic! i am still not getting the correct answer... the algebra is killing me.

OpenStudy (anonymous):

not sure myself, I'm not getting that answer...

OpenStudy (anonymous):

you sure that's right?

OpenStudy (anonymous):

looks like the partial rather than implicit diff...

OpenStudy (anonymous):

the answer is supposed to be \[-\frac{ xy }{ x^2+2y^2 }\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

that's correct... didn't you put xy/(sqrt(x^2+y^2)) above?

OpenStudy (anonymous):

i did im sorry, that was incorrect so i deleted it to avoid confusion

OpenStudy (anonymous):

\[y'\sqrt{x^2+y^2} + \frac{ y^2y' }{ \sqrt{x^2+y^2} } = \frac{ -xy }{\sqrt{x^2+y^2} }\]

OpenStudy (anonymous):

had an extra ' 2' in there...

OpenStudy (anonymous):

\[ \frac{ y'(x^2+y^2)+y^2y' }{ \sqrt{x^2+y^2} } = \frac{ -xy }{\sqrt{x^2+y^2} }\]

OpenStudy (anonymous):

idk where it comes from

OpenStudy (anonymous):

\[ y' = \frac{ -xy }{\sqrt{x^2+y^2} }\frac{ \sqrt{x^2+y^2} }{ (x^2+y^2)+y^2 }\]

OpenStudy (anonymous):

\[ y' = \frac{ -xy }{1 }\frac{ 1 }{ (x^2+y^2)+y^2 }\]

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

sure:)